OpenStudy (nsh4267):

4 3/8 divided by 1 1/2 reduced to lowest terms

Hero:

You are given: \(\dfrac{4 + \dfrac{3}{8}}{1 + \dfrac{1}{2}}\) Multiply top and bottom by LCD/GCF \(\dfrac{\left(4 + \dfrac{3}{8}\right)}{\left(1 + \dfrac{1}{2}\right)} \times \dfrac{\dfrac{1}{8}}{\dfrac{1}{8}}\) Distribute to get \(\dfrac{32 + 3}{8 + 4}\) Which simplifies to: \(\dfrac{35}{12}\) Perform the following steps to write as mixed fraction: \(\dfrac{24 + 11}{12}\) \(\dfrac{24}{12} + \dfrac{11}{12}\) \(2 + \dfrac{11}{12}\)